OpenStudy (anonymous):

Anybody good at Algebra 3?

OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m = -1/2 and the point (-2, -3).

OpenStudy (anonymous):

im not but I can guess you the wrong answer

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

standard form = ax+by = c first put the equation in point slope form y-y1=m(x-x1) m = the slope then put it into standard form

OpenStudy (anonymous):

ax + by = c. How do I figure out what the letters are? @helpme1.2

OpenStudy (anonymous):

like i said, first you need to input it in point slope form then it will get easier y+3 = -1/2(x+3)--distrubute the -1/2 y+3 = -1/2x-3/2 -- subtract 3 to both sides y = -1/2x-9/2 now for last step put y = -1/2x-9/2 into the standard form

OpenStudy (mathmale):

We are given m=-1/2 and the point (-2,-3). The point-slope form of the equation of this line is thus y-(-3) = (-1/2)(x-[-2]), or y+3 = (-1/2)(x+2). Clear out the fractions by multiplying both sides of the equation by 2: 2y+6=-(x+2) = -x - 2. Add +x to both sides. Subtract 6 from both sides. Simplify.

OpenStudy (mathmale):

Is this still the problem on which you wish to focus?

OpenStudy (anonymous):

Yes! I just can't seem to get it... :|

OpenStudy (mathmale):

Please look at my previous comments; can you finish the work as I asked?

OpenStudy (anonymous):

x + 2y = -8.